OpenStudy (anonymous):

how do you write an equation of the line passing through (-3,2) and is perpendicular to the line y=6x-2?

OpenStudy (anonymous):

y=6x-2 m1 = 6 Since lines are perpendicular m1*m2 = -1 6 * m2 = -1 m2 = -1/6 Nw using point Slope Form.... (-3,2) y-y1 = m2(x-x1).....Hope nw u can Solve..